OpenStudy (anonymous):

Charlie was carrying around the steps to solve Square root of 6x plus 4 = 8. On his way through the Math Factory, Charlie dropped them and they mixed with other steps. Put the correct steps in order to solve Square root of 6x plus 4 = 8. A. Divide by 2 from both sides B. Square root both sides C. Divide by 6 from both sides D. Subtract 4 from both sides E. Subtract 8 from both sides F. Square both sides

OpenStudy (anonymous):

@Imyourbulbasaur

OpenStudy (anonymous):

@CoolRanchDoritos

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

@rational

OpenStudy (anonymous):

lol we can do this

OpenStudy (anonymous):

\[\sqrt{6x+4}=8\] right ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

lets to it ans write the steps \[\sqrt{6x+4}=8\] 1) square both sides \[6x+4=64\]

OpenStudy (anonymous):

2) subtract \(4\) from both sides \[6x=60\]

OpenStudy (anonymous):

3) divide both sides by \(6\) \[x=10\]

OpenStudy (anonymous):

that is all the rest are not steps

OpenStudy (anonymous):

now i have a question

OpenStudy (anonymous):

wait so its b

OpenStudy (anonymous):

lol you must be confused the answer is not a letter you are told " Put the correct steps in order "

OpenStudy (anonymous):

wow i was dumb lost thank you

OpenStudy (anonymous):

you got it now?